OpenStudy (anonymous):

Solve the system by the addition method: 3x-4y=-7 2x-7y=4 I cant find the solution so I dont think there is one, let me know if you can find it please.

OpenStudy (anonymous):

Just multiply the first equation by 2 and the second one by -3, summing them will yield \[13y=-26\] so y=-2. Replacing y in one of the equation gives x=-5.

OpenStudy (anonymous):

ya but you do that for the other equation and you get -25/3

OpenStudy (anonymous):

shouldnt they have the same result all across the board?

OpenStudy (anonymous):

I didn't understand your point, checking the result for both equations is ok...

OpenStudy (anonymous):

okay so we solved for y and y=-2 so yes for 2x-7y=4, you would get the result that x=-5 , but if you plug it into the other equation which is 3x-4y=-7 the result for x is x= -25/3 . if y=-2 then shouldnt x have a consistent value for all of the equations?

OpenStudy (anonymous):

\[3x-4(-2)=-7\]\[3x+8=-7\]\[3x=-8-7\]\[3x=-15\]\[x=-5\]right?

OpenStudy (anonymous):

yup. huge mathematical error on my part haha

OpenStudy (anonymous):

i skipped a step and put my first equation as 3x+8=-17

OpenStudy (anonymous):

instead of -7, dumbbb thank you

OpenStudy (anonymous):

One confession: first time I did a mistake also and ended up with x=-1/3...

OpenStudy (anonymous):

ya I guess were rushing haha

OpenStudy (anonymous):

just to be sure, the solution is (-5,-2) correct?